.According to a survey by the Administrative Management Society, one-half of U.S. companies give employees 4 weeks of vacation after they have been with the company for 15 years. Find the probability that among G companies surveyed at random, the number that give employees 4 weeks of vacation after 15 years of employment is (a) anywhere from 2 to 5; (b) fewer than 3.
1
Expert's answer
2019-12-10T11:05:22-0500

Let "X=" the number of the companies that give employees 4 weeks of vacation after 15 years of employment: "X\sim B(n;p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x},x=0, 1, 2,...,n"

Given that "n=G, p=0.5"

(a) Find the probability that among G companies surveyed at random, the number that give employees 4 weeks of vacation after 15 years of employment is anywhere from 2 to 5


"P(2\leq X\leq5)=P(X=2)+P(X=3)+""+P(X=4)+P(X=5)=""=\binom{G}{2}0.5^2(1-0.5)^{G-2}+\binom{G}{3}0.5^3(1-0.5)^{G-3}+""+\binom{G}{4}0.5^4(1-0.5)^{G-4}+\binom{G}{5}0.5^4(1-0.5)^{G-4}=""=0.5^G\bigg({G! \over 2!(G-2)!}+{G! \over 3!(G-3)!}+{G! \over 4!(G-4)!}+{G! \over 5!(G-5)!}\bigg),""G\geq5"

If "G=6"


"P(2\leq X\leq5)=""=0.5^6\bigg({6! \over 2!(6-2)!}+{6! \over 3!(6-3)!}+{6! \over 4!(6-4)!}+{6! \over 5!(6-5)!}\bigg)=""=0.875"

(b) Find the probability that among G companies surveyed at random, the number that give employees 4 weeks of vacation after 15 years of employment is fewer than 3.


"P(X<3)=P(X=0)+P(X=1)+P(X=2)=""=\binom{G}{0}0.5^0(1-0.5)^{G-0}+\binom{G}{1}0.5^1(1-0.5)^{G-1}+""+\binom{G}{2}0.5^2(1-0.5)^{G-2}=""=0.5^G\bigg({G! \over 0!(G-0)!}+{G! \over 1!(G-1)!}+{G! \over 2!(G-2)!}\bigg),x\geq2"

If "G=6"


"P(X<3)=0.5^6\bigg({6! \over 0!(6-0)!}+{6! \over 1!(6-1)!}+{6! \over 2!(6-2)!}\bigg)=""=0.34375"